Mr. Lee brought reading books to his class. If each student takes 5 books. there will be 15 books left. If 5 students do not take a book and the rest of the students take 7 books each, there will be no books left. How many books were brought to the class?
A) 140 B) 125 C) 120 D) 104

Found 2 solutions by Boreal, zjjc123:
Answer by Boreal(15235)

x books were brought to class.
y=number of students in the class.
x-5y=15
x-7(y-5)=0. This means that 5 fewer than y students each took 7 books.
Therefore, x=7(y-5)
Substitute into first equation.
7(y-5)-5y=15
7y-35-5y=15
2y=50
y=25
x-125=15
x=140 books. A
There are 25 students.
If each student takes 5 books, 125 are taken and 15 are left
If 20 students (5 fewer than 25) take 7 books, none is left.

Answer by zjjc123(1)

y=5x+15 y=7(x-5)
y=5x+15 y=7x-35
5x+15=7x-35
15=2x-35
2x=50
x=25
25 students
140 books